# Quench dynamics of the quantum XXZ chain with staggered interactions: Exact results and simulations on digital quantum computers

### v2

@@ Abstract
We investigate quench dynamics in the quantum $S=1/2$ XXZ antiferromagnetic chain with staggered and anisotropic interactions in the flat-band limit. Our quench protocol interchanges the odd- and even-bond strengths of a fully dimerized chain, enabling us to derive exact time-dependent states for arbitrary even system sizes by working in the Bell basis. We obtain closed-form, size-independent expressions for the von Neumann and second-order Rényi entanglement entropies. We further calculate exact Loschmidt echoes and the corresponding return rate functions across various anisotropies and system sizes, and identify Loschmidt zeros in finite chains. Our analysis reveals the precise conditions on the anisotropy parameter that govern the periodicity of the dynamical observables. In addition to the analytic study, we perform two types of numerical experiments on IBM-Q quantum devices. First, we use the Hadamard test to estimate the Bell-basis expansion coefficients and reconstruct the dynamical states, achieving accurate entanglement entropies and the Loschmidt echo for small systems. Second, we implement Trotter-error-free time-evolution circuits combined with randomized Pauli measurements. Post-processing via statistical correlations and classical shadows yields reliable estimates of the second-order Rényi entanglement entropy and the Loschmidt echo, showing satisfactory agreement with exact results.
@@ I Introduction
In this paper, we investigate quench dynamics of the quantum $S=1/2$ XXZ antiferromagnetic chain with staggered and anisotropic interactions in the flat-band limit. We consider global quenches implemented by interchanging the odd- and even-bond strengths in a fully dimerized XXZ chain. This system fails to relax at long times after the quench due to the vanishing group velocity of its excitations. We derive a formula for the time-dependent states in the Bell basis for arbitrary finite system size, and obtain closed-form solutions for dynamical entanglement entropies quantified by both the von Neumann and the second-order Rényi entropies. We also compute the exact Loschmidt echo, identify Loschmidt zeros in certain finite chains, and determine analytical expressions for a variety of anisotropy parameters and system sizes. From these exact results, we establish the conditions on the combined parameter $J\Delta$, the product of bond strength $J$ and anisotropy $\Delta$, that determine the periodicity or non-periodicity of the dynamical observables.
@@ II Model and its non-equilibrium dynamics
For $\Delta=0$, the model reduces to the XX chain, which can be mapped to well-studied free-fermion systems [45, 46]. At $\Delta=1$, the system becomes the XXX chain: in the uniform case ($J=J^{\prime}$), it is exactly solvable via the Bethe ansatz [47], while in the dimerized case it exhibits the well-known spin-Peierls transition [48, 49].
@@ II.2 Dynamical quantities
The entanglement entropy (EE) quantifies the degree of quantum entanglement between a subsystem and its complement. For a time-evolved state $\left|{\Psi_{0}(t)}\right\rangle$ partitioned into two parts $A$ and $B$, the entanglement between them is encoded in the reduced density matrix of one part, say $A$,
@@
where $\text{Tr}_{B}$ denotes the partial trace over $B$. We consider both the von Neumann EE, defined by
@@
and the second-order Rényi EE, given by
@@
For the quench protocols studied here, both the von Neumann and Rényi EE exhibit persistent oscillations, indicating the absence of relaxation due to confinement [30, 31]. Such persistent entanglement oscillations have been previously reported in the fully dimerized XX chain with $\Delta=0$ (equivalently, the flat-band Su-Schrieffer-Heeger chain) [52, 13]. Here we extend this analysis to the XXZ chain with $\Delta\neq 0$.
@@
In the limit of large $N$, the Loschmidt echo has an asymptotic form given by
@@ III.1 Half-chain entanglement entropy
The key ingredient of the entanglement entropy is the reduced density matrix, $\rho_{A}$, of the subsystem considered. Here for a half chain, the dimension of the reduced density matrix is $2^{n}\times 2^{n}$. Using the eigenvalues of $\rho_{A}$, that are all real and satify $0\leq\lambda_{l}(t)\leq 1,\;\sum_{l=1}^{2^{n}}\lambda_{l}(t)=1$, we obtain the von Neumann EE by
@@ III.1.1 $n=2$
Thus, the von Neumann EE can be obtained by inserting the eigenvalues into Eq. (21), yielding
@@
Similarly, the Rényi EE is obtained from Eq. (22):
@@
For the $X$-chain with $\Delta=0$, the expression for the von Neumann EE is reduced to
@@ III.1.2 $n=3,5,\cdots$
To clarify this, Fig. 2 illustrates the initial singlet-product state (gray lines) and the couplings (red lines) in the postquench Hamiltonian for chains of length $n=2,4$ (Fig. 2(a) and (b)) and $n=3,5$ (Fig. 2 (c) and (d)), where the dashed line indicates the symmetric bipartition. In our setup, each initial entangled dimer located at $(i,i+1)$ can spread over three lattice spacings, covering sites from i1i-1 to $i+2$, during the time evolution [63, 13]. The EE of a half chain is the number of entangled states that connect sites inside the partition to sites outside. For any odd $n>2$, the initial state always contains an entangled dimer that crosses the boundary between sites $i=n$ and $i=n+1$, contributing a time-independent unit to the EE; this is the constant term ’$1$’ in Eq. (33). In addition, at most two entangled states may cross the boundary between sites $i=1$ and $i=2n$, evolved from the singlets initially at $(1,2)$ and (2n1,2n)(2n-1,2n). This time-dependent part of the EE is analogous to the $n=2$ case. However, in the $n=2$ chain, the short chain length forces a single entangled state to cross both boundaries simultaneously, whereas for $n>2$ it can cross only one boundary. As a result, one unit of the EE in the $n=2$ case involves two couplings with a total strength $2J$ in the postquench Hamiltonian (red lines in Fig. 2), while for $n>2$ it involves only a single postquench coupling of strength $J$. This difference leads to a doubling of the entanglement oscillation period for odd $n>2$ compared with the $n=2$ case.
@@ III.1.3 $n=4,6,\cdots$
Repeating the above reasoning, we can also find an exact relation between the dynamical EE of even $n>2$ and that of $n=2$. Similar to the $n=2$ case, no singlet-dimer initially crosses the bipartition boundaries. The dynamical EE arises solely from two pairs of dimers initially located adjacent to the boundaries, one pair at (n,n1)(n,n-1) and $(n+1,n+2)$, and the other at $(1,2)$ and (2n1,2n)(2n-1,2n). For $n>2$, each pair of dimers is connected through one postquench coupling of strength $J$, and each dimer can cross only one boundary during the time evolution. In contrast, in a $n=2$ chain there exists only one such pair coupled through two postquench couplings ($2J$), and each entangled state crosses the two boundaries at the same time. Therefore, the dynamical EE for even $n>2$ is related to that of the $n=2$ case by
@@
From Eq. (34), the von Neumann EE for an even $n>2$ chain is explicitly given by
@@
For $n=4$, the analytical forms of the eigenvalues of the $16$ reduced density matrices for the XX- and XXX-chains are provided in Appendix A.
@@ III.2 Loschmidt echo
Figure 4: Time dependence of the exact Loschmidt echo for the XXZ chain of length N=2n,n=2,4,6N=2n,\,n=2,4,6 and $8$ with $J=1$ and various values of the anisotropy parameter: (a) $\Delta=0$, corresponding to the XX chain; (b) $\Delta=1/2$; (c) $\Delta=1$, the XXX chain; (d) $\Delta=7/4$.
Figure 5: Time dependence of the return rate function calculated from the Loschmidt echo in Fig. 4 for the XXZ chain with $J=1$ and the anisotropy parameter at (a) $\Delta=0$, corresponding to the XX chain; (b) $\Delta=1/2$; (c) $\Delta=1$, the XXX chain; (d) $\Delta=7/4$.
Figure 6: Time dependence of the exact Loschmidt echo for the XXZ chain with odd $n=3,5,7$ and at various values of the anisotropy parameter: (a) $\Delta=0$, corresponding to the XX chain; (b) $\Delta=1/2$; (c) $\Delta=1$, the XXX chain; (d) $\Delta=7/4$. All curved are plotted with fixed $J=1$.
Figure 7: Time dependence of the return rate function calculated from the Loschmidt echo in Fig. 6 for the XXZ chain with $J=1$ and the anisotropy parameter at (a) $\Delta=0$, corresponding to the XX chain; (b) $\Delta=1/2$; (c) $\Delta=1$, the XXX chain; (d) $\Delta=7/4$.
@@
Below we present closed-form solutions of short XXZ chains with $n=2,3,4$ and XX chains up to $n=8$. For the shortest chain with $n=2$, the dimer configurations for nonzero coefficients are given in Eq. (23), leading to the Loschmidt echo
@@
$\left.+\sin^{2}(J\Delta t)\sin^{4}(Jt)\right\}\;,$ (41)
which for the special cases $\Delta=0$ (XX chain) is reduced to
${\cal L}_{XX}^{n=4}(t)=\left[\frac{1}{4}\left(1+\cos^{2}(Jt/2)\right)^{2}\right]^{2}\;.$ (42)
and for $\Delta=1$ (XXX chain) is
${\cal L}_{XXX}^{n=4}(t)=\frac{1}{16}\left(1-12\cos^{2}(Jt/2)+42\cos^{4}(Jt/2)\right$.
$\left.-36\cos^{6}(Jt/2)+21\cos^{8}(Jt/2)\right)\;.$ (43)
For larger systems, we provide the analytical forms of ${\cal L}(t)$ for the $n=6$ $X$-chain:
${\cal L}_{XX}^{n=6}(t)=\left(\frac{17}{32}\cos^{2}(Jt/2)+\frac{15}{32}\cos^{4}(Jt/2)\right)^{2}\;,$ (44)
and for the $n=8$ $X$-chain:
${\cal L}_{XX}^{n=8}(t)=\left\{\left[(1+7\cos^{2}(Jt/2))/8\right]^{2}-\frac{1}{4096}\sin^{2}(2Jt)\right$.
$\left.-\frac{7}{1024}\left[5\sin^{2}(Jt)+\cos(Jt)-\cos(3Jt)\right]\right\}^{2}\;.$ (45)
The closed-form solution of the Loschmidt echo for $n=6$ with an arbitrary $\Delta$ is given in Appendix B.
@@
Moreover, we observe exact zeros of the Loschmidt echo for even $n$. For instance, the XX chain ($\Delta=0$) with $n=2$ has Loschmidt zeros at $t=\pi,3\pi,\cdots$; the $n=4$ XXZ chain with $\Delta=1/2$, and also the longer $n=8$ chain with $\Delta=7/4$ exhibit zeros at $t=\pi,3\pi,\cdots$, too. At these instances, the return rate function becomes divergent, as shown in Fig. 5. In all cases considered here, Loschmidt zeros do not appear in odd $n$ chains [Fig. 6 and 7], nor in the XXX chain for any $n$. For the irrational value $\Delta=2/(\sqrt{5}+1)$, we find no Loschmidt zeros up to $t=8\pi$ [see Fig. 8 (b) and (d)].

### v3

@@ Abstract
We investigate quench dynamics in the quantum $S=1/2$ XXZ antiferromagnetic chain with staggered and anisotropic interactions in the flat-band limit. Our quench protocol interchanges the odd- and even-bond strengths of a fully dimerized chain, enabling us to derive exact time-dependent states for arbitrary even system sizes by working in the Bell basis. We obtain closed-form, size-independent expressions for the von Neumann and second-order Rényi entanglement entropies. We further calculate exact Loschmidt echoes and the corresponding return rate functions across various anisotropies and system sizes, and identify Loschmidt zeros in finite chains. Our analysis reveals distinct finite-size scaling of the Loschmidt echo at critical times with chain length and identifies the precise conditions on the anisotropy parameter governing the periodicity of the dynamical observables. In addition to the analytic study, we perform two types of numerical experiments on IBM-Q quantum devices. First, we use the Hadamard test to estimate the Bell-basis expansion coefficients and reconstruct the dynamical states, achieving accurate entanglement entropies and the Loschmidt echo for small systems. Second, we implement Trotter-error-free time-evolution circuits combined with randomized Pauli measurements. Post-processing via statistical correlations and classical shadows yields reliable estimates of the second-order Rényi entanglement entropy and the Loschmidt echo, showing satisfactory agreement with exact results.
@@ I Introduction
In this paper, we investigate quench dynamics of the quantum $S=1/2$ XXZ antiferromagnetic chain with staggered and anisotropic interactions in the flat-band limit. We consider global quenches implemented by interchanging the odd- and even-bond strengths in a fully dimerized XXZ chain. This system fails to relax at long times after the quench due to the vanishing group velocity of its excitations. We derive a formula for the time-dependent states in the Bell basis for arbitrary finite system size, and obtain closed-form solutions for dynamical entanglement entropies quantified by both the von Neumann and the second-order Rényi entropies. We also compute the exact Loschmidt echo, identify Loschmidt zeros in certain finite chains, and determine analytical expressions for a variety of anisotropy parameters and system sizes. From these exact results, we establish the conditions on the combined parameter $J\Delta$, the product of bond strength $J$ and anisotropy $\Delta$, that determine the periodicity or non-periodicity of the dynamical observables. Furthermore, we uncover special finite-size scaling behaviors of the Loschmidt echo at dynamical phase transitions, which, to the best of our knowledge, have not been reported previously.
@@ II Model and its non-equilibrium dynamics
For $\Delta=0$, the model reduces to the $X$ chain, which can be mapped to well-studied free-fermion systems [45, 46]. At $\Delta=1$, the system becomes the XXX chain: in the uniform case ($J=J^{\prime}$), it is exactly solvable via the Bethe ansatz [47], while in the dimerized case it exhibits the well-known spin-Peierls transition [48, 49].
@@ II.2 Dynamical quantities
The entanglement entropy ($E$) quantifies the degree of quantum entanglement between a subsystem and its complement. For a time-evolved state $\left|{\Psi_{0}(t)}\right\rangle$ partitioned into two parts AA and $B$, the entanglement between them is encoded in the reduced density matrix of one part, say AA,
@@
where $\text{Tr}_{B}$ denotes the partial trace over $B$. We consider both the von Neumann $E$, defined by
@@
and the second-order Rényi $E$, given by
@@
For the quench protocols studied here, both the von Neumann and Rényi $E$ exhibit persistent oscillations, indicating the absence of relaxation due to confinement [30, 31]. Such persistent entanglement oscillations have been previously reported in the fully dimerized $X$ chain with $\Delta=0$ (equivalently, the flat-band Su-Schrieffer-Heeger chain) [52, 13]. Here we extend this analysis to the XXZ chain with $\Delta\neq 0$.
@@
In the limit of large NN, the Loschmidt echo has an asymptotic form given by
@@ III.1 Half-chain entanglement entropy
The key ingredient of the entanglement entropy is the reduced density matrix, $\rho_{A}$, of the subsystem considered. Here for a half chain, the dimension of the reduced density matrix is $2^{n}\times 2^{n}$. Using the eigenvalues of $\rho_{A}$, that are all real and satify $0\leq\lambda_{l}(t)\leq 1,\;\sum_{l=1}^{2^{n}}\lambda_{l}(t)=1$, we obtain the von Neumann $E$ by
@@ III.1.1 $n=2$
Thus, the von Neumann $E$ can be obtained by inserting the eigenvalues into Eq. (21), yielding
@@
Similarly, the Rényi $E$ is obtained from Eq. (22):
@@
Both functions ${\cal S}_{1}^{n=2}(t)$ and ${\cal S}_{2}^{n=2}(t)$ reach their maximum value of 22, independent of $\Delta$, at $t=(2m+1)\pi/2J$ with $m=0,1,2,\cdots$ when the condition $\cos(Jt)=0$ is satisfied.
For the $X$-chain with $\Delta=0$, the expression for the von Neumann $E$ is reduced to
@@ III.1.2 $n=3,5,\cdots$
To clarify this, Fig. 2 illustrates the initial singlet-product state (gray lines) and the couplings (red lines) in the postquench Hamiltonian for chains of length $n=2,4$ (Fig. 2(a) and (b)) and $n=3,5$ (Fig. 2 (c) and (d)), where the dashed line indicates the symmetric bipartition. In our setup, each initial entangled dimer located at $(i,i+1)$ can spread over three lattice spacings, covering sites from i1i-1 to $i+2$, during the time evolution [63, 13]. The $E$ of a half chain is the number of entangled states that connect sites inside the partition to sites outside. For any odd $n>2$, the initial state always contains an entangled dimer that crosses the boundary between sites $i=n$ and $i=n+1$, contributing a time-independent unit to the $E$; this is the constant term ’11’ in Eq. (33). In addition, at most two entangled states may cross the boundary between sites $i=1$ and $i=2n$, evolved from the singlets initially at $(1,2)$ and (2n1,2n)(2n-1,2n). This time-dependent part of the $E$ is analogous to the $n=2$ case. However, in the $n=2$ chain, the short chain length forces a single entangled state to cross both boundaries simultaneously, whereas for $n>2$ it can cross only one boundary. As a result, one unit of the $E$ in the $n=2$ case involves two couplings with a total strength $2J$ in the postquench Hamiltonian (red lines in Fig. 2), while for $n>2$ it involves only a single postquench coupling of strength $J$. This difference leads to a doubling of the entanglement oscillation period for odd $n>2$ compared with the $n=2$ case.
@@ III.1.3 $n=4,6,\cdots$
Repeating the above reasoning, we can also find an exact relation between the dynamical $E$ of even $n>2$ and that of $n=2$. Similar to the $n=2$ case, no singlet-dimer initially crosses the bipartition boundaries. The dynamical $E$ arises solely from two pairs of dimers initially located adjacent to the boundaries, one pair at (n,n1)(n,n-1) and $(n+1,n+2)$, and the other at $(1,2)$ and (2n1,2n)(2n-1,2n). For $n>2$, each pair of dimers is connected through one postquench coupling of strength $J$, and each dimer can cross only one boundary during the time evolution. In contrast, in a $n=2$ chain there exists only one such pair coupled through two postquench couplings ($2J$), and each entangled state crosses the two boundaries at the same time. Therefore, the dynamical $E$ for even $n>2$ is related to that of the $n=2$ case by
@@
From Eq. (34), the von Neumann $E$ for an even $n>2$ chain is explicitly given by
@@
For $n=4$, the analytical forms of the eigenvalues of the $16$ reduced density matrices for the $X$- and XXX-chains are provided in Appendix A.
@@ III.2 Loschmidt echo
Figure 4: Time dependence of the exact Loschmidt echo for the XXZ chain of length N=2n,n=2,4,6N=2n,\,n=2,4,6 and 88 with $J=1$ and various values of the anisotropy parameter: (a) $\Delta=0$, corresponding to the $X$ chain; (b) $\Delta=1/2$; (c) $\Delta=1$, the XXX chain; (d) $\Delta=7/4$.
Figure 5: Time dependence of the return rate function calculated from the Loschmidt echo in Fig. 4 for the XXZ chain with $J=1$ and the anisotropy parameter at (a) $\Delta=0$, corresponding to the $X$ chain; (b) $\Delta=1/2$; (c) $\Delta=1$, the XXX chain; (d) $\Delta=7/4$.
Figure 6: Time dependence of the exact Loschmidt echo for the XXZ chain with odd $n=3,5,7,9$ and at various values of the anisotropy parameter: (a) $\Delta=0$, corresponding to the $X$ chain; (b) $\Delta=1/2$; (c) $\Delta=1$, the XXX chain; (d) $\Delta=7/4$. All curved are plotted with fixed $J=1$.
Figure 7: Time dependence of the return rate function calculated from the Loschmidt echo in Fig. 6 for the XXZ chain with $J=1$ and the anisotropy parameter at (a) $\Delta=0$, corresponding to the $X$ chain; (b) $\Delta=1/2$; (c) $\Delta=1$, the XXX chain; (d) $\Delta=7/4$.
@@
Below we present closed-form solutions of short XXZ chains with $n=2,3,4,6$ and $X$ chains up to $n=8$. For the shortest chain with $n=2$, the dimer configurations for nonzero coefficients are given in Eq. (23), leading to the Loschmidt echo
@@
$\left.+\sin^{2}(J\Delta t)\sin^{4}(Jt)\right\}\;.$ (41)
An analytical expression for $n=6$ with arbitrary values of $\Delta$ is given in Appendix B.
For the special case $\Delta=0$ ($X$ chain), we have
${\cal L}_{XX}^{n=2}(t)=\frac{1}{4}$ $\left[1+\cos(Jt)\right]^{2}\;,$ (42)
${\cal L}_{XX}^{n=4}(t)=\frac{1}{4^{5}}$ $\left[19+12\cos(Jt)+\cos(2Jt)\right]^{2}\;,$ (43)
${\cal L}_{XX}^{n=6}(t)=\frac{1}{4^{9}}$ $\left[226+255\cos(Jt))+30\cos(2Jt)\right$.
$\left.+\cos(3Jt)\right]^{2}\;,$ (44)
${\cal L}_{XX}^{n=8}(t)=\frac{1}{4^{13}}$ $\left[3225+3976\cos(Jt)+924\cos(2Jt)\right$.
$\left.+56\cos(3Jt)+\cos(4Jt)\right]^{2}\;,$ (45)
where the Loschmidt echo for even $n$ is written as a quadratic form in a superposition of $\cos(\mu Jt)$-terms with $\mu=0,1,\cdots,n/2$ and a global prefactor $C_{n}=1/4^{2n-3}$. Simialrly, for odd $n$, it can be written as linear combination of $\cos(\mu Jt)$-terms with $\mu=0,1,\cdots,n-1$ and $C_{n}=1/4^{2n-2}$. Explicit expressions for odd $n=5,7$ and 99 are provided in Appendix C. Extensions to larger $n$ are straightforward. These expressions also imply that the corresponding Loschmidt amplitude, $\langle{\Psi_{0}}|{\Psi_{0}(t)}\rangle$, is real for even $n$ and is complex for odd $n$.
@@
Moreover, we observe exact zeros of the Loschmidt echo for even $n$. For instance, the $X$ chain ($\Delta=0$) with $n=2$ has Loschmidt zeros at $t=\pi,3\pi,\cdots$; the $n=4$ XXZ chain with $\Delta=1/2$, and also the longer $n=8$ chain with $\Delta=7/4$ exhibit zeros at $t=\pi,3\pi,\cdots$, too. At these instances, the return rate function becomes divergent, as shown in Fig. 5. In all cases considered here, Loschmidt zeros do not appear in odd $n$ chains [Fig. 6 and 7], nor in the XXX chain for any $n$. For the irrational value $\Delta=2/(\sqrt{5}+1)$, we find no Loschmidt zeros up to $t=8\pi$ [see Fig. 8 (b) and (d)].
III.3 Dynamical quantum phase transitions and finite-size scaling
Since the prequench and postquench Hamiltonians belong to distinct topological phases, the quench considered here is expected to exhibit dynamical phase transitions [64, 59].
Under a fermionic transformation, the $X$ chain decouples into two transverse-field Ising chains [65, 66]. Within this mapping, the present protocol by assuming $J=1$ then corresponds to a quench from a fully polarized initial state along the transverse-field direction, $\left|{\Psi_{0}}\right\rangle=\bigotimes_{i}\left|{\rightarrow_{i}}\right\rangle$, evolved under a classical Ising Hamiltonian. Equivalently, it can be viewed as a quench from a fully polarized state into a cluster phase [67] governed by ${\cal H}_{\text{cluster}}=-1/4\sum_{i}Z_{i-1}X_{i}Z_{i+1}$, where $X$ and $Z$ denote Pauli matrices. Previous studies [68] have established that such quenches exhibit dynamical quantum phase transitions at critical times $t^{*}=(2m+1)\pi$, with $m=0,1,2,\cdots$.
Figure 9: Visualization of the Loschmidt echo for the XXZ chain with odd $n=5,7,9$ and various values of $\Delta$, where the label $\Delta=0.618$ indicates an irrational value $\Delta=2/(\sqrt{5}+1)$. At $t=\pi$, all the curves intersect exactly at $(1/2)^{2(n-1)}$ (dashed line). The coupling strength is fixed at $J=1$.
[…]
Case (i)
For $n=4\nu+2,\nu=0,1,2,\cdots$:
${\cal L}^{*}_{XX}=0,\;r^{*}=\infty\,.$
Case (ii)
For $n=4\nu+4,\nu=0,1,2,\cdots$:
${\cal L}^{*}_{XX}=(1/2)^{2(n-2)},\;r^{*}=(1-2/n)\ln 2\,.$
Case (iii)
For odd $n$:
${\cal L}^{*}_{XX}=(1/2)^{2(n-1)},\;r^{*}=(1-1/n)\ln 2\,.$
[…]
These results indicate that quenches in the flat-band limit of the XXZ chain exhibit dynamical phase transitions at $t^{*}=(2m+1)\pi$, independent of $\Delta$, where $\ln{\cal L}_{XXZ}(t^{*})\to\infty$ and $r(t^{*})\to\ln 2$ in the thermodynamic limit. For certain finite system sizes with even $n$, additional exact zeros of ${\cal L}_{XXZ}$ occur at times $t\neq t^{*}$, for example at even multiples of $\pi$ [see Fig. 5].
A further noteworthy feature is that the dynamical phase transitions considered here are also characterized by maximum entanglement entropy for $n>2$. At the critical times $t^{*}=(2m+1)\pi$ with $J=1$, the entanglement entropy reaches ${\cal S}=4$ for even $n>2$ and ${\cal S}=3$ for odd $n$, independent of $\Delta$, as can be verified from Eqs. (33) and (34).